OpenStudy (anonymous):

Find the exact value of the expression: log 2 base 16. I did the following: 16^x = 2 2^4x=2 4x=2 x=1/2

OpenStudy (mathmate):

16^x=2 (definition of logarithm) 2^(4x) = 2^(1) (both sides raised to a power of base 2) Can you complete your work from here?

OpenStudy (kohai):

Are you allowed to use a calculator?

OpenStudy (solomonzelman):

\(\large\color{slate}{\log_22^4 }\) .....

OpenStudy (solomonzelman):

4 goes on the outside.

OpenStudy (solomonzelman):

and on

OpenStudy (anonymous):

Got 1/4 as my answer. Found my mistake. Thanks for the help!

OpenStudy (mathmate):

Great, 1/4 is good! :)